Greenhab Report – December 3rd

rew 200 GreenHab Report – 03-DEC-2018

Greenhab Officer: Makiah Eustice

Environmental Control:


Shade Cloth on

Average Temperatures:

Low: 70 F

High: 80 F

Hours of Supplemental Light: 5

Daily Water Usage of Crops: 10 L

Water in Blue Tank (lbs): 1/3 full

Times of Watering for Crops: 1000

Changes to crops:

Planted 3 small cups of Water Cress for Andrew’s school outreach. Planted shallow bed of sun flower seeds and shallow bed of mixed salad greens. All reported in the Greenhab Log


First day in the Greenhab! A sauna for me, but a retreat for anyone trying to briefly escape the cold. Still trying to understand the operations and needs of individual plants.

Harvest: NONE

Support/Supplies Needed: The humidity/temperature measure is not displaying the temperature correctly; we may need a new sensor. The supplemental light did not come on Sunday night, but did tonight; I will check to see if the light goes off at midnight tonight. Some of the snap peas and tomatoes are not doing well; I believe they need more sun to thrive.

Sol Summary – December 3rd

Health and Safety Officer Report
Lindsay Rutter, Health and Safety Officer of Crew 200

Crew 200 Sol Summary Report 03Dec2018

Sol: 1
Summary Title: Introduced guests and fixed solar observatory Author’s name: Lindsay Rutter
Mission Status: All crew members are healthy and productive. Sol Activity Summary: We interacted with guests and fixed solar observatory. Look Ahead Plan: We will do science most of the day tomorrow. Anomalies in work: Solar observatory had issues that needed fixing. Weather: Unexpected snowy weather created muddy and slippery ground. Crew Physical Status: One crew member felt headache due to cold weather, rested, and has no more symptoms.
EVA: No EVAs today due to inclement weather.
Reports to be file:
Support Requested:

Operations Report – December 02nd

Crew 200 Operations Report 02-Dec-2018


Name of person filing report: Shannon Rupert

Non-nominal systems: generator autostart

Notes on non-nominal systems: nothing to report

Generator (hours run): From 5:30 pm last night to 2 am this afternoon (due to snow). On tonight at 4:30 pm.

Solar— SOC 91 %

Diesel Reading – not noted %

Propane Reading – 49 %

Ethanol Free Gasoline – 4 gallons

Water (auxiliary tank) – 0 gallons

Water (static tank) – 500 gallons

Auxiliary to Static tank transfer– no

Gallons transferred: 0

Water in GreenHab – not noted gallons, getting low

Water (loft) – Static to Loft Pump used – no

Water Meter: not recorded

Toilet tank emptied: no

Deimos rover used: no, still not functional


Beginning charge:

Ending charge:

Currently charging: no

Sojourner rover used: yes


Beginning charge: 100

Ending charge:

Currently charging: yes

Spirit rover used: yes


Beginning charge: 74

Ending charge:

Currently charging: yes

Opportunity rover used: yes


Beginning charge: 85?

Ending charge:

Currently charging: yes

Curiosity rover used: yes


Beginning charge: 100

Ending charge:

Currently charging: yes

Notes on rovers: Two rovers parked in front of the RAM at new charging stations. Spirit and Opportunity should be parked there, while Curiosity should be parked next to the GreenHab until Deimos is operational. We are still seeing much shorter ranges on the rovers, not sure if it is a problem or just lack of use, so we will have to give them more ride time to see if this improves.

# Hours the ATVs were Used today: less than one

Notes on ATVs: ATVs were taken out of active service and are now parked away from the station

HabCar used and why, where? No

CrewCar used and why, where? Yes, just moved to its regular parking space

General notes and comments: The crew worked amazingly hard yesterday repairing the tunnel and getting it covered. We managed to get it all up and functional, although there is still much work to be done to tidy up the tarps and secure it for the season. But this is work that can be done in sim or by a single person, so the day was a complete success.

When I shut down the power for the installations of the RAM’s power, I had to shut down the entire system, which I didn’t know how to do. So I did it the only way I know how and now the charger button is acting differently than before, but it still fixes the problem of generator speed so I guess we are fine until the contractor arrives next week.

Summary of internet: Nominal.

Summary of suits and radios: The crew used the suits and so can give a complete report on their condition as soon as possible.

Summary of Hab operations: Rodent found a new home.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: The crew helped to rearrange the bench modules to reflect the new design of the space. The ovens/incubator are now on the metal bench. The heater will be installed on the lower bench of this table and should not be moved from that spot.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing at this time.